# LaT-PFN: A Joint Embedding Predictive Architecture for In-context Time-series Forecasting

### v1

@@ 4 Experimental Setup and Training Details
During training, we employ a linear warmup schedule for weight decay, and target encoder ema decay – in line with previous JEPA work [5, 6]. We trained the final model for 24 hours on a single NVIDIA A10G Tensor Core GPU – per seed. Additionally, we employ tuning on schedule parameters and $\mu$-parametrization, which was essential for training stability [59]. For more training and tuning details, see Appendix C.
@@
We report the mean and error statistics of the zero-shot performance of multiple seeds. These are accuracy, MSE per time-step, and the cumulative relative root mean squared error (CRRMSE). The latter is normalized by the cumulative target. We motivate choosing CRRMSE for being scale-invariant, and practical applicability in tasks such as sales forecasting.
@@
For real datasets, fixed sequence length contexts were provided to the model. These were mapped to an abstract interval and Z-normalized with 2-std using the history. Which samples to include as contexts is coined Context Curation, and is a process analogous to Prompt Engineering in Large Language Models (LLMs) [40, 56]. For details, see Appendix C.
@@
We selected FBProphet [51], ARIMA [10], and ForecastPFN [17] as forecasting baselines, all of which have open-source codebases. We compared against TS2Vec, in a downstream classification task, [63] to evaluate the quality of the embedding space. We use the original authors’ codebase and either download the trained model weights or re-train, wherever applicable. Finally, all baselines are tuned, where applicable, and then evaluated on identical datasets. For more details on baselines, see Appendix C.
Table 1: Forecasting scores (with std) for different time horizons. Results demonstrate a strong performance for all of the benchmarks from decoded latent predictions. For details see Appendix C
@@ 5 Results & Discussion
Table 2: Overall accuracy of UCR datasets, by fitting SVM on top of fixed-length summary embeddings [20]. Per-dataset scores are in Appendix E.
training budget accuracy
TS2Vec finetuned 45.7%
LaT-PFN zero-shot 51.1%
@@
Next, we investigate the importance of the context by evaluating the uplift in performance on a synthetic validation set, for different context sizes. We chose to carry out this experiment on simulated data to guarantee a control environment. The result is visualized in Figure 8, which shows a clear correlation between performance and context size. More interestingly, this pattern persists even beyond the original context size of the model during training (represented by the vertical dotted line). This validates our assumption that the key to carrying out in-context learning for zero-shot forecasting is to provide a rich and substantial set of examples. This can be especially impactful when applied in a production setting with an abundance of example series, such as sales forecasting.
@@
We evaluate the quality of LaT-PFN’s embedding space by introducing a second downstream task, namely time series classification on the 129 multi-domain UCR datasets. We compare our model to TS2Vec, a state-of-the-art universal representation learning framework for time series [63]. We train TS2Vec on the UCR datasets, then fit Support Vector Machines (SVM) [8] on their embedding space. We compare this to an SVM fitted on the fixed-length summary embeddings of a frozen LaT-PFN, trained on synthetic data only. As is evident in the results reported in Table 2, LaT-PFN outperforms TS2Vec across the entire benchmark of datasets. See Appendix E for individual dataset scores.
We argue these results may be due to bias introduced by time series-specific heuristics and contrastive positive pair selection, both of which TS2Vec and similar methods rely heavily upon [63, 60]. By contrast, LaT-PFN independently discovers the most representative embedding space in which self-supervised learning is most optimal. Furthermore, TS2Vec is trained on the datasets it is evaluated on. This may traditionally be thought of as an advantage. Yet, often these datasets are small and lack variety, increasing the risk of out-of-distribution data in the test set. On the other hand, LatPFN’s prior ensures a wider variety of contexts, whilst adapting to the specific context during test time.
@@
We carry out a qualitative analysis of the latent space generated by LaT-PFN. When analyzing the fixed-length summary embeddings after dimensionality reduction, we observe a clear separation in clusters by dataset – see Figure 7. This suggests that LaT-PFN picks up distributional differences between these domains, making these fixed-length summary embeddings useful for search, classification and other downstream applications.
Figure 9: Embeddings per-timestep show patch-like behavior. More in Appendix D
Interestingly we can observe the emergence of distinct, regular, and discrete patch-like tokens in the heatmaps of the embedding space. An example of this is visualized in Figure 9. Whilst not always humanly interpretable, discernible patterns often emerge from a closer analysis of these patches. For instance, Figure 10 plots some examples in the data space, alongside their closest matches when taking the $L_{2}$-distance between the average patch-embedding. We can observe that these are visually related and seem to describe specific local features. This presents similarities to the patch-like processing of visual "tokens" in Vision Transformers (ViT) and their approach to high-level feature representation [18]. LaT-PFN, however, appears to have learned these tokens independently, without the need to outright encode this pattern in the model’s architecture.
Finally, after applying dimensionality reduction to the set of patch embeddings (Figure 11), we notice the clear emergence of a few clusters. Given the aforementioned similarity between shapes of patches in data space with similar token embeddings, we speculate these could be a latent vocabulary that has been independently learned by the model, consisting of common "words" amongst multi-domain series. This would once more validate our core assumption that, despite time series data often being perceived as highly domain-dependant, there is in fact enough commonality to warrant the use of meta-learning approaches to create foundational models for this data modality.
@@ 6 Related Work
For time series forecasting, focus has recently shifted from statistical models such as ARIMA [10], ES [25] and FBProphet [51] towards deep learning. Early pioneers are DeepAR [47] and hybrid methods [49]. Subsequently, N-Beats [42], N-hits [12], Informer [65] and FEDformer [66], combine classical with deep learning paradigms.
More recently, following the success of the NLP domain, many attempts have been made to leverage the transformer architecture [54] to create foundation models [9] for time series forecasting. The two main approaches for doing so can be categorized as; (a) re-purposing LLMs [26, 67, 31] and (b) training transformer models from scratch [17, 4, 57, 15, 61, 45]. By contrast, our research seeks to integrate representation- and in-context learning in time series forecasting, combining strong performance with a low computational budget.
Another approach altogether is to create embeddings and leave the downstream application to the discretion of end users. Notable are TS2Vec [63], TimeCLR [60], SimMTM [16] and TF-C [64]. These methods often rely heavily on expert heuristics, whereas our approach serves both as an embedder and a forecaster, instead leveraging a self-supervised latent space to drive predictions.
@@ 7 Limitations
The research presented has a few limitations. Firstly, the current framework is constrained to univariate series, limiting its applicability to multivariate scenarios. Similarly, this model has yet to prove effective in handling hierarchical, discrete, or zero-inflated series, which are often underrepresented in this research field. Defining appropriate synthetic priors could address this, which we leave to future work. Secondly, we observe a low level of standardization for time series deep learning research. Whilst we found an (over) abundance of standard datasets, we observed that data processing, and target and horizon selection are still arbitrary, reducing the comparability of results across studies. This is beyond the scope of this paper but may be addressed in future work. Finally, we designed our model to be highly adaptive to a given context. As a result, sub-optimal contexts can lead to variability in performance. We argue this is analogous to prompt engineering, which assigns this responsibility to the expert user. In future work, we hope to address this by exploring automated methods – via retrieval augmented generation (RAG) [37] or prompt tuning [36]. Addressing these points will be valuable for enhancing the model’s versatility.
@@ 8 Conclusion
In this paper, we proposed LaT-PFN, a novel architecture for zero-shot univariate time series forecasting. The model integrates both the PFN and JEPA architecture to allow for in-context learning in latent space. LaT-PFN is trained exclusively on synthetic data, for which we define a novel prior. Combined with an abstract time dimension, this makes our model extremely versatile and able to achieve strong forecasting results across different data distributions, time granularities, and forecast horizons. Finally, its strong embedding space exhibits interesting emergent qualities, such as the development of patch-like tokens that resemble a corpus of time series features. This makes LaT-PFN ideal for supporting transfer learning on downstream models.
@@
Appendix A Broader Societal Impacts
@@
Appendix B Derivation PFN NLLLoss
This derivation shows how to approximate Bayesian Inference via a PFN. It is adapted from [41] and forms the justification of why using a simulation with a cross-entropy loss is an approximation of the PPD. Assume we want to approximate PPD $P(y|x,\mathcal{D})$ with network $Q(y|x,\mathcal{D})$. We start with the KL-divergence between the two and end up at entropy function $\bm{H}()$ and constant $C$. The latter does not depend on parameters $\theta$ and is dropped in optimisation.
@@ Appendix B Derivation PFN NLLLoss
We can then integrate the PFN simulation by sampling the data from our prior:
@@
Appendix C Reproducibility Details
@@
C.1 Exhaustive Description Synthesis Training Data
@@
Unlike [17], we train using an abstract time dimension. In order to include many different types of temporal patterns, we define a resolution parameter $\rho$, which determines how many points are represented per interval, for each individual time series. We define this as the relation between the sequence dimension and resolution parameter:
@@ C.2 Baselines
  • ForecastPFN [17] is a flexible in-context forcaster model and one of the inspirations for this work. We download their open-source repository available under Apache Licence 2.0 and trained model weights. We adopt their inference script to work in our repository. We have tried several data normalizations by hand to get the most out their model and picked the best one, which turned out to be Z-normalization with 2std – on top of the normalizations present in the repository already.
@@ C.3 Benchmark Preprocessing
For each dataset, we create univariate time series out of each of the columns. Additionally, we further split up the univariate series along the time dimension to create single entities of different granularity, which we then use for context creation. We normalize the time dimension of each dataset so that it fits in the abstract $[-3,1]$ interval. Furthermore, we apply Z-score normalization, with 2 standard deviations, to the values of each univariate time series.
@@ C.3.1 Illness
We hand-picked 4 windows of the historic series as context examples and kept 1 as held-out starting from forecast date 01-01-2018, which we evaluate the zero-shot performance on.
@@
C.3.2 EttH1 & EttH2
@@
    ETTh2 OT 2017/01 - 2017/05
C.3.3 Traffic
@@ C.3.4 ECL
We provide a context dimension of 15 and leave two held-out time series.
@@
OT 2019/01/01 - 2018/01/14 OT 2019/05/01 - 2018/05/14
OT 2019/01/07 - 2018/01/20 OT 2019/04/01 - 2018/04/14
OT 2019/01/17 - 2018/01/30
OT 2019/02/01 - 2019/02/14
OT 2019/02/07 - 2019/02/20
OT 2019/02/17 - 2019/03/02
OT 2019/03/01 - 2019/03/14
OT 2019/03/07 - 2019/03/20
OT 2019/03/17 - 2019/03/30
OT 2017/04/01 - 2017/04/14
OT 2017/04/07 - 2017/04/20
OT 2017/04/17 - 2017/04/30
C.4 Training & Tuning
@@
  • We tune all of the above hyperparameters with Optuna [3] in the 3 runs, spanning a total of 216 hours.
@@
  • We apply TF32 quantization but no mixed-precision training since this gave stability issues
@@
  • We map our training data to an abstract time domain of $[-3,1]$, with the forecast moment at the origin.
  • We define 100 bins on the abstract $[-3.5,3.5]$ range, as the output space for the decoder.
@@
Appendix D Latent Space Visualised

### v2

@@ 4 Experimental Setup and Training Details
During training, we employ a linear warmup schedule for weight decay, and target encoder ema decay – in line with previous JEPA work [4, 5]. We trained the final model for 24 hours on a single NVIDIA A10G Tensor Core GPU – per seed. Additionally, we employ tuning on schedule parameters and $\mu$-parametrization, which was essential for training stability [59]. For more training and tuning details, see Appendix D.
@@
We report the mean and error statistics, assumed normally distributed, of the zero-shot performance of five seeds. These are accuracy, MSE per time-step, and the cumulative relative root mean squared error (CRRMSE). The latter is normalized by the cumulative target. We motivate CRRMSE for its scale-invariance and practical applicability in tasks such as sales forecasting.
@@
For real datasets, contexts were divided in fixed windows, normalized to a fixed time interval, and Z-normalized with 2-std using the history. The context-sample selection process is coined Context Curation, and is analogous to Prompt Engineering in Large Language Models (LLMs) [40, 56]. For dataset-specific details see Appendix D.
@@
We selected FBProphet [51], ARIMA [9], and ForecastPFN [16] as forecasting baselines, all of which have open-source codebases. We compared against TS2Vec, in a downstream classification task, [63] to evaluate the quality of the embedding space. We use the original authors’ codebase and either download the trained model weights or re-train, wherever applicable. Finally, all baselines are tuned, where applicable, and then evaluated on identical datasets. For more details on baselines, see Appendix D.
Table 1: Forecasting scores (with std) for different time horizons. Results demonstrate a strong performance for all of the benchmarks from decoded latent predictions. See data details in Appendix D
@@ 5 Results & Discussion
Table 2: Overall accuracy (with std) of all UCR datasets, by fitting SVM on top of fixed-length summary embeddings [19]. Per-dataset scores are in Appendix F.
train budget accuracy
TS2Vec 35 epochs 44.8% $\pm$ 1.1%
LaT-PFN zero-shot 47.9% $\pm$ 3.7%
@@
Next, we investigate the importance of the context by evaluating the uplift in performance on a synthetic validation set, for different context sizes. We chose to carry out this experiment on simulated data to guarantee a controlled environment. The result is visualized in Figure 8, which shows a clear correlation between performance and context size. More interestingly, this pattern persists even beyond the original context size of the model during training (represented by the vertical dotted line). This validates our assumption that the key to carrying out in-context learning for zero-shot forecasting is to provide a rich and substantial set of examples. This can be especially impactful when applied in a production setting with an abundance of example series, such as sales forecasting.
@@
We evaluate the quality of LaT-PFN’s embedding space by introducing a second downstream task, namely time series classification on the 128 multi-domain UCR datasets [19]. We compare our model to TS2Vec, a state-of-the-art universal representation learning framework for time series [63]. We train TS2Vec on the UCR datasets, then fit Support Vector Machines (SVM) [7] on their embedding space. We compare this to an SVM fitted on the fixed-length summary embeddings of a frozen LaT-PFN, trained on synthetic data only. As is evident in the results reported in Table 2, LaT-PFN outperforms TS2Vec on the combined datasets, despite the former being run in the zero-shot setting. See Appendix F for individual dataset scores.
We argue these results may be due to bias introduced by time series-specific heuristics and contrastive positive pair selection, both of which TS2Vec and similar methods rely heavily upon [63, 60]. By contrast, LaT-PFN independently discovers a representative embedding space in which self-supervised learning is highly efficient. Furthermore, TS2Vec is trained on the datasets it is evaluated on. This may traditionally be thought of as an advantage. Yet, often these datasets are small and lack variety, increasing the risk of out-of-distribution data in the test set. On the other hand, LatPFN’s prior ensures a wider variety of contexts, whilst adapting to the specific context during test time.
@@
We carry out a qualitative analysis of the latent space generated by LaT-PFN. When analyzing the fixed-length summary embeddings after dimensionality reduction, we observe a clear separation in clusters by dataset – see Figure 7. This suggests that LaT-PFN picks up distributional differences between these domains, making the fixed-length summary embeddings useful for search, classification, and other downstream applications.
Figure 9: Normalized embeddings across sequence show patch-like behavior. More in Appendix E
Interestingly, in the per-timestep embeddings we can observe the emergence of distinct, regular, and discrete patch-like tokens in the heatmaps of the embedding space. An example of this is visualized in Figure 9. Whilst not always humanly interpretable, discernible patterns often emerge from a closer analysis of these patches. For instance, Figure 10 plots some examples in the data space, alongside their closest matches when taking the $L_{2}$-distance between the average patch-embedding. We can observe that these are visually related and seem to describe specific local features. This presents similarities to the patch-like processing of visual "tokens" in Vision Transformers (ViT) and their approach to high-level feature representation [17]. LaT-PFN, however, appears to have learned these tokens independently, without the need to outright encode this pattern in the model’s architecture.
Finally, after applying dimensionality reduction to the set of patch embeddings (Figure 11), we notice the clear emergence of a few clusters. Considering the noted similarity between the shapes of patches in data space, and their corresponding averaged patch "token" embeddings, we hypothesize that these patches may represent a latent vocabulary independently learned by the model, comprising common "words" across multi-domain series. This would once more validate our core assumption that, despite time series data often being perceived as highly domain-dependant, there is in fact enough commonality to warrant the use of meta-learning approaches to create foundational models for this data modality.
@@ 6 Related Work
For time series forecasting, the focus has recently shifted from statistical models such as ARIMA [9], ES [25] and FBProphet [51] towards deep learning. Early pioneers are DeepAR [47] and hybrid methods [49]. Subsequently, N-Beats [42], N-hits [11], Informer [65] and FEDformer [66], combine classical with deep learning paradigms.
Following the success of the NLP domain, many attempts have been made to leverage the transformer architecture [54] to create foundation models [8] for time series forecasting. The two main approaches for doing so can be categorized as; (a) re-purposing LLMs [26, 67, 31] and (b) training transformer models from scratch [16, 3, 57, 14, 61, 45]. By contrast, our research seeks to integrate representation- and in-context learning in time series forecasting, combining strong performance with a low computational and data budget.
Another approach altogether is to create embeddings and leave the downstream application to the discretion of end users. Notable are TS2Vec [63], TimeCLR [60], SimMTM [15] and TF-C [64]. These methods often rely heavily on expert heuristics, whereas our approach serves both as an embedder and a forecaster, leveraging a self-supervised latent space to drive predictions.
@@ 7 Limitations
The research presented has a few limitations. Firstly, the current framework is constrained to univariate series, limiting its applicability to multivariate scenarios. Similarly, this model has yet to prove effective in handling hierarchical, discrete, or zero-inflated series, which are often underrepresented in this research field. Defining appropriate synthetic priors could address this, which we leave to future work. Secondly, we observe a low level of standardization for time series deep learning research. Whilst we found an (over) abundance of standard datasets, we observed that data processing, and target and horizon selection are still arbitrary, reducing the comparability of results across studies. This is beyond the scope of this paper but may be addressed in future work. Thirdly, we designed our model to be highly adaptive to a given context. As a result, sub-optimal contexts can lead to variability in performance. We argue this is analogous to prompt engineering, which assigns this responsibility to the expert user. In future work, we hope to address this by exploring automated methods – via retrieval augmented generation (RAG) [37] or prompt tuning [36]. Finally, the model has proven somewhat sensitive to initialization and the normalization functions used. Addressing these points will be valuable for enhancing the model’s versatility.
@@ 8 Conclusion
In this paper, we proposed LaT-PFN, a novel architecture for zero-shot univariate time series forecasting. The model integrates both the PFN and JEPA architecture to allow for in-context learning in latent space. LaT-PFN is trained exclusively on synthetic data, for which we define a novel prior. Combined with a normalized abstract time-axis, this makes our model extremely versatile and able to achieve strong forecasting results across different data distributions, time granularities, and forecast horizons. Finally, its strong embedding space exhibits interesting emergent qualities, such as the development of patch-like tokens that resemble a corpus of time series features. This makes LaT-PFN ideal for supporting transfer learning on downstream models.
@@
Appendix A Code
We provide the code for this work at https://github.com/StijnVerdenius/Lat-PFN.
Appendix B Broader Societal Impacts
@@
Appendix C Derivation PFN NLLLoss
This derivation, adapted from [41], illustrates the use of a cross-entropy loss with a synthetic prior as an approximation of the Predictive Posterior Distribution (PPD). This is a more effective way to learn Bayesian inference, as opposed to estimating the true posterior, which is often intractable. Furthermore, it yields a flexible approach by encoding expert data, which makes it highly adaptable for developing downstream applications [41].
Assume we want to approximate PPD $P(y|x,\mathcal{D})$ with network $Q(y|x,\mathcal{D})$, for contexts $\mathcal{D}$. We start with the expectation over the contexts, of the KL-divergence between distributions $P(.)$ and $Q(.)$:
@@ Appendix C Derivation PFN NLLLoss
We find equivalence to the entropy function $\bm{H}(.)$ and constant $C$. The latter in independent of parameters $\theta$ and is dropped in optimization [41]. Next, we elect to incorporate the PFN synthetic prior simulation by sampling data from our simulation:
@@
These formulas show that, for the purpose of optimization, this is equivalent to applying the cross-entropy loss – a commonly used loss function. The output space is split in bins to allow the use of a cross-entropy loss for regression and forecasting problems – following the PFN example [41]. Furthermore, the synthetic prior provides control over the prior distribution of contexts and access to infinite training data resources, including contexts, held-out examples, and targets [41].
Finally, in this work, we introduce a separation of concerns between predicting and decoding. The end-to-end PPD task is effectively split into (a) learning a good way of summarizing the prior distribution within the latent prediction distribution $P(\hat{\bar{y}}|x,\mathcal{D},\theta)$, and then (b) separately learning an approximated posterior based on that summary $P(y|\hat{\bar{y}},\theta)$ – again without assumptions on output distribution family. Although the optimization is decoupled, there remains a causal relationship between the output distribution over $y$ and it dependent variables: $x$, $\mathcal{D}$, and parameters $\theta$. Therefore, we argue that this approach is a reasonable approximation of the PPD – given the latent predictions as an intermediary.
Appendix D Reproducibility Details
@@
D.1 Exhaustive Description Synthesis Training Data
@@
Unlike [16], we train using an normalized time dimension. In order to include many different types of temporal patterns, we define a resolution parameter $\rho$, which determines how many points are represented per interval, for each individual time series. We define this as the relation between the sequence dimension and resolution parameter:
@@ D.2 Baselines
  • ForecastPFN [16] is a flexible in-context forecasting model and one of the inspirations for this work. We download their open-source repository available under Apache Licence 2.0 and trained model weights. We adopt their inference script to work in our repository. We have tried several data normalizations by hand to get the most out their model and picked the best one, which turned out to be Z-normalization with 2std – on top of the normalizations present in the repository already.
@@ D.3 Benchmark Preprocessing
For each dataset, we create univariate time series out of each of the columns. Additionally, we further split up the univariate series along the time dimension to create single entities of different granularity, which we then use for context creation. We normalize the time dimension of each dataset so that it fits within the normalized $[-3,1]$ interval. Furthermore, we apply Z-score normalization, with 2 standard deviations, to the values of each univariate time series.
@@ D.3.1 Illness
We hand-picked 8 windows of the historic series as context examples and kept 1 as held-out starting from forecast date 01-01-2018, which we evaluate the zero-shot performance on.
@@
The illness dataset is provided by a public government institution (CDC), which does not provide an explicit license.
D.3.2 EttH1 & EttH2
@@
    ETTh2 HULL 2017/01 - 2017/05
Both datasets are under Creative Commons Attribution-NoDerivatives 4.0 International.
D.3.3 Traffic
@@ D.3.4 ECL
We provide a context dimension of 30 and leave two held-out time series.
@@
208 2018/04/01 - 2018/04/14 OT 2019/05/01 - 2018/05/14
208 2018/04/07 - 2018/04/20 OT 2019/04/01 - 2018/04/14
208 2018/04/17 - 2018/04/30
208 2017/04/01 - 2017/04/14
208 2017/04/07 - 2017/04/20
208 2017/04/17 - 2017/04/30
208 2019/01/01 - 2019/01/14
208 2019/01/07 - 2019/01/20
208 2019/01/17 - 2019/01/30
208 2019/02/01 - 2019/02/14
208 2019/02/07 - 2019/02/20
208 2019/02/17 - 2019/03/02
208 2019/03/01 - 2019/03/14
208 2019/03/07 - 2019/03/20
208 2019/03/17 - 2019/03/30
313 2018/04/01 - 2018/04/14
313 2018/04/07 - 2018/04/20
313 2018/04/17 - 2018/04/30
313 2017/04/01 - 2017/04/14
313 2017/04/07 - 2017/04/20
313 2017/04/17 - 2017/04/30
313 2019/01/01 - 2019/01/14
313 2019/01/07 - 2019/01/20
313 2019/01/17 - 2019/01/30
313 2019/02/01 - 2019/02/14
313 2019/02/07 - 2019/02/20
313 2019/02/17 - 2019/03/02
313 2019/03/01 - 2019/03/14
313 2019/03/07 - 2019/03/20
313 2019/03/17 - 2019/03/30
This dataset is under CC BY 4.0 licence.
D.4 Training & Tuning
@@
  • We tune all of the above optimizer hyperparameters, $\lambda_{\text{di}}$ and $\lambda_{\text{latent}}$ with Optuna [2] in the 3 runs, spanning a total of 216 hours.
@@
  • We apply TF32 quantization but no mixed-precision training since this gave stability issues.
@@
  • We map our training data to an normalized time domain of $[-3,1]$, with the forecast moment at the origin.
  • We define 100 bins on the normalized $[-3.5,3.5]$ range, as the output space for the decoder.
@@
Appendix E Latent Space Visualised